# Conservative Perception Models for Probabilistic Verification

### v2

@@ 4. Problem Formulation: Conservative Verification
where $x(t)\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state (with bounded $\mathcal{X}$); $y(t)\in\mathbb{R}^{p}$ are the observations; $z(t)\in\{0,1\}$ is the output of a perception system [^1] For clarity of presentation, we assume the perception system detects the presence of a single obstacle in the environment. However, our approach can handle systems with multiple types of obstacles and $z(t)$ from a fixed finite set., $u(t)\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control output, the functions $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$, $g:\mathbb{R}^{n}\times\mathbb{R}^{v}\to\mathbb{R}^{p}$, $h:\mathbb{R}^{p}\to\mathbb{R}^{z}$ describe the system dynamics, sensor maps, and perception maps; the function $c:\mathbb{R}^{p}\to\mathbb{R}^{m}$ is a stateless controller; and $v(t)\in D_{v}\subseteq\mathbb{R}^{v}$ describes perception noise. The $v(t)$ noise represents inexact sensing, such as a LiDAR scanner giving noisy distance estimates.
@@
Given system $\mathsf{M}_{\text{sys}}$ as in (1), a model MDP or IMDP $\mathsf{M}_{\text{abs}}$, and safety property $\psi$, we say the $\mathsf{M}_{\text{abs}}$ is a conservative system model with respect to $\mathsf{M}_{\text{sys}}$ if
$P^{min}_{\mathsf{M}_{\text{abs}}}(\psi)\leq P_{\mathsf{M}_{\text{sys}}}(\psi)$
@@
Therefore, the controller-plant model $\mathsf{M}_{\text{cpl}}$ can be constructed as a conservative non-probabilistic MDP (i.e., one whose behavior contains that of true system $\mathsf{M}_{\text{sys}}$) with a standard discrete abstraction technique (for example, in (cleaveland_monotonic_2022, )). However, constructing a conservative $\textsf{M}_{\text{per}}$, either as an MDP or IMDP, is an open problem. Therefore, this paper focuses on synthesizing $\textsf{M}_{\text{per}}$ in a conservative manner using labeled perception data (i.e., the perception outputs $y$ paired with the true underlying states $x$) from the system $\mathsf{M}_{\text{sys}}$. In Section 6, we will show that our $\mathsf{M}_{\text{abs}}$ is a conservative system model, as per Definition 1.
@@
$\mathsf{M}_{\text{abs}}=\mathsf{M}_{\text{cpl}}~{}||~{}\textsf{M}_{\text{per}}$
@@
Our controller-plant model $\mathsf{M}_{\text{cpl}}$ is a non-probabilistic MDP with $\mathsf{M}_{\text{cpl}}=\{S_{\mathsf{M}_{\text{cpl}}},s^{0}_{\mathsf{M}_{\text{cpl}}},\alpha_{\mathsf{M}_{\text{cpl}}},\delta_{\mathsf{M}_{\text{cpl}}},L_{\mathsf{M}_{\text{cpl}}}\}$. The (discrete) state space $S$ is formed by partitioning the (continuous) state space $\mathcal{X}$ of $\mathsf{M}_{\text{sys}}$ into a finite set of mutually exclusive regions, so every $s\in S_{\mathsf{M}_{\text{cpl}}}$ has a corresponding region $X\subseteq\mathcal{X}$ (we use $x\in s$ to denote this correspondence). The initial state $s^{0}_{\mathsf{M}_{\text{cpl}}}$ is the state in $S_{\mathsf{M}_{\text{cpl}}}$ which contains the initial state $x_{0}$ of $\mathsf{M}_{\text{sys}}$ in its set of corresponding states. The actions $\alpha_{\mathsf{M}_{\text{cpl}}}=\alpha_{per}\times\alpha_{reach}$, where $\alpha_{per}=\{0,1\}$ (which we refer to as “perception actions”) are the outputs of the perception component $h$ and $\alpha_{reach}$ (which we refer to as “reachability actions”) are the actions that correspond to the non-determinism required to conservatively model the continuous system dynamics $f$ and controller $u$ with a discrete abstraction. [^2] For a visualization of this phenomenon, see Figure 2 in (cleaveland_monotonic_2022, ). The transition relation $\delta_{\mathsf{M}_{\text{cpl}}}$ maps states and a pair (perception action, reachability actions) to the next state in $S_{\mathsf{M}_{\text{cpl}}}$. Finally, $L_{\mathsf{M}_{\text{cpl}}}:S\rightarrow 2^{AP}$, were $AP$ is a set of atomic propositions.
@@
$\mathsf{M}_{\text{cpl}}=\{S_{\mathsf{M}_{\text{cpl}}},s_{0},\alpha_{\mathsf{M}_{\text{cpl}}},\delta_{\mathsf{M}_{\text{cpl}}},L_{\mathsf{M}_{\text{cpl}}}\}$ is a conservative controller/plant abstraction if for every state $s\in S_{\mathsf{M}_{\text{cpl}}}$ and perception action $a_{per}\in\alpha_{per}$, the set of next states $S^{\prime}=\{s^{\prime}\in S|\exists a_{reach}\in\alpha_{reach}\text{ s.t. }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{a_{per},a_{reach}\},s^{\prime})=1\}$ in $\mathsf{M}_{\text{cpl}}$ contains every state in $\mathsf{M}_{\text{sys}}$ that can be mapped to by the dynamics of $\mathsf{M}_{\text{sys}}$. Formally:
$\forall s\in S_{\mathsf{M}_{\text{cpl}}},\forall a_{per}\in\alpha_{per},\forall x\in s,\exists s^{\prime}\in S,\exists a_{reach}\in\alpha_{reach}\in\delta_{\mathsf{M}_{\text{cpl}}}(s,a)$
such that $\text{such that }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{a_{per},a_{reach}\},s^{\prime})=1\text{ and }f(x,u(a))\in s^{\prime}$
@@
With the above in mind, this paper solves the following problem:
@@
Given the closed-loop dynamical system $\mathsf{M}_{\text{sys}}$ defined in Equation 1, a safety property $\psi$, perception data $\mathcal{D}_{\text{sys}}=\{x_{i},z_{i}\}_{i=1,\ldots,N}$ where $x_{i}\in\mathbb{R}^{n}$ is the state of $\mathsf{M}_{\text{sys}}$ and $z_{i}\in\{0,1\}$ is the output of the perception component, and a conservative model $\mathsf{M}_{\text{cpl}}$ of the control and dynamics portion of $\mathsf{M}_{\text{sys}}$, synthesize a model $\textsf{M}_{\text{per}}$ of the perception component(s) of $\mathsf{M}_{\text{sys}}$ such that
$P^{min}_{\mathsf{M}_{\text{cpl}}||\textsf{M}_{\text{per}}}(\psi)\leq P_{\mathsf{M}_{\text{sys}}}(\psi)$
5. Conservative Approach to Perception Modeling
@@
    For each discrete state, compute an interval of detection probabilities using $\mathcal{D}$ and binomial confidence intervals (fleiss_statistical_2004, ). This is a novel step where we differ from existing perception models (badithela_evaluation_2023, ; cleaveland_conservative_2023, ).
@@
Figure 1. Overview of our approach to conservative perception modeling.
5.1. Step 1: State Space Discretization
@@
In terms of performing the binning in practice, we see two main ways of doing this. First, one can divide the state space into a set of equally spaced hyper-rectangles (or intervals if the state space is 1D, rectangles in 2D, …) [^3] This is the approach taken by the previous works (badithela_evaluation_2023, ; cleaveland_conservative_2023, ).. In this case, the size of the hyper-rectangle serves as a hyper-parameter of the approach. The second way is to first select the desired number of bins for each dimension of the state space $\mathcal{X}$, which we denote as $\mathcal{b_{1}},\ldots,\mathcal{b_{n}}$, and then divide the state space such that each bin has an equal number of datapoints in it. For example, if one has a 1D system and $l$ datapoints (so $|\mathcal{D}|=l$), then one would sort the points $\{x_{i},z_{i}\}\in\mathcal{D}$ by the $x_{i}$ values and assign each chunk of $\frac{\mathcal{l}}{\mathcal{b}}$ to its own bin.
@@
We point out two other, subtler trade-offs at play that our discretization-based approach identifies. The existing perception abstraction approaches (badithela_evaluation_2023, ; cleaveland_conservative_2023, ) fail to account for these trade-offs, since they ignore the inherent uncertainty present in $\textsf{M}_{\text{per}}$— and so they fail to make it conservative. The first tradeoff is that using more bins means that there will be fewer datapoints per bin. This will lead to larger probability intervals for each bin in Step 2 since there are fewer datapoints fed into the binomial confidence intervals. These larger detection probability intervals in turn will lead to a larger interval on the estimated safety probability $\hat{P}(\psi)$. This trade-off captures the intuition that having fewer datapoints to compute the transition probabilities of $\textsf{M}_{\text{per}}$ generally results in a less precise model.
@@
Figure 2. Larger discretization bins lead to smaller binomial confidence intervals (blue) but necessitate more probability enlargement (orange).
5.2. Step 2: Bin Probability Computation
@@
$\mathcal{D}_{s}=\{(x,z)\in\mathcal{D}~{}|~{}x\in s\}$
Just as in prior works (badithela_evaluation_2023, ; cleaveland_conservative_2023, ), we can then compute the empirical detection probability of the datapoints in $\mathcal{D}_{s}$:
(2) $\hat{p}_{s}=\frac{\sum_{(x,z)\in\mathcal{D}_{s}}\mathbbm{1}_{z=1}}{|\mathcal{D}_{s}|}$
@@
To alleviate this issue, we replace the point estimate $\hat{p}_{s}$ with a confidence interval (CI), denoted as $[\underline{\hat{p}}_{s},\bar{\hat{p}}_{s}]$ for bin $s$ [^4] Using CIs to construct IMDPs has been explored in the areas of synthesis (jiang_safe_2022, ) and active learning (alasmari_quantitative_2022, ) (which inspires our approach), but not verification. There are several methods to obtain a confidence interval for the detection probability of the perception model within the bin, also known as a binomial confidence interval. We employ the Clopper-Pearson confidence interval because its coverage level is never less than the desired nominal coverage (clopper_use_1934, ). When constructing a confidence interval, one needs to select a confidence level $1-\alpha$. Using a larger value of $1-\alpha$ leads to larger probability intervals, but the proportion of bins in which the confidence interval contains the true detection probability will, in turn, be larger.
We argue that the choice of $1-\alpha$ should be determined by the desired confidence in the model checking result for the overall system $\mathsf{M}_{\text{sys}}$. If one wants a confidence of $1-\alpha_{MC}$ that the model checking results are conservative (i.e.  Definition 1 holds with probability $1-\alpha_{MC}$) and there are $|S|$ states in $\textsf{M}_{\text{per}}$, then one should set $1-\alpha=1-\alpha_{MC}/|S|$. Intuitively, this formula captures the fact that each bin’s confidence interval has an $\alpha_{MC}/|S|$ probability of not containing the true detection probability, so by applying a union bound (boole_mathematical_1847, ), the probability of every bin’s confidence interval containing the true detection probability is lower bounded by
(3) $1-\sum_{s\in S}\frac{\alpha_{MC}}{|S|}=1-\alpha_{MC}$
@@
5.3. Step 3: Probability Enlargement
@@
(4) $\delta^{p}_{s}:=\left(\max_{x\in s}\;LR(x)\right)-\left(\min_{x\in s}\;LR(x)\right)$
@@
(5) $[\underline{\hat{p}}_{s}-\delta^{p}_{s},\bar{\hat{p}}_{s}+\delta^{p}_{s}]$
@@
If we fit a surrogate model to detection probabilities, why not use it to construct $\textsf{M}_{\text{per}}$? We attempted to directly use the logistic regression, along with its associated confidence interval, to estimate the range of true detection probabilities within the bins. However, we found that the logistic regression confidence interval did not empirically achieve the coverage that we wanted. That is because the logistic regression CIs can be viewed as the CIs for the parameters of how to map data onto the space of logistic models — not for the data itself. We demonstrate this in the experiments in Section 7.4 and particularly Figure 5(a).
6. Theoretical Guarantee of Conservative Verification
In this section, we state the main theoretical guarantee of this work. Specifically, we show that given the inputs to Problem 1 and an $\textsf{M}_{\text{per}}$ constructed using Steps 1–3 from Section 5, verifying the resulting composed model $\mathsf{M}_{\text{abs}}$ will give a conservative estimate of the safety of the true system $\mathsf{M}_{\text{sys}}$:
(6) $P_{\mathsf{M}_{\text{sys}}}(\psi)\in[P_{\mathsf{M}_{\text{abs}}}^{min}(\psi),P_{\mathsf{M}_{\text{abs}}}^{max}(\psi)]$

### v3

@@ IV Problem: Conservative Verification
where $x(t)\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state (with bounded $\mathcal{X}$); $y(t)\in\mathbb{R}^{p}$ are the observations; $z(t)\in\{0,1\}$ is the output of a perception system [^1] For clarity of presentation, we assume the perception system detects the presence of a single obstacle in the environment. However, our approach can handle systems with multiple types of obstacles and $z(t)$ from a fixed finite set., $u(t)\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control output, the functions $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$, $g:\mathbb{R}^{n}\times\mathbb{R}^{v}\to\mathbb{R}^{p}$, $h:\mathbb{R}^{p}\to\{0,1\}$ describe the system dynamics, sensor maps, and perception maps; the function $c:\{0,1\}\to\mathbb{R}^{m}$ is a stateless controller; and $v(t)\in D_{v}\subseteq\mathbb{R}^{\mathcal{v}}$ describes (potentially unbounded) perception noise. The $v(t)$ noise represents inexact sensing, such as a LiDAR scanner giving noisy distance estimates. Due to the noise, Equation 7 is a stochastic dynamical system.
@@
Given system $\mathsf{M}_{\text{sys}}$ as in Equation 7, a model MDP or IMDP $\mathsf{M}_{\text{abs}}$, and safety property $\psi$, we say the $\mathsf{M}_{\text{abs}}$ is a conservative system model with respect to $\mathsf{M}_{\text{sys}}$ if
$P^{min}_{\mathsf{M}_{\text{abs}}}(\psi)\leq P_{\mathsf{M}_{\text{sys}}}(\psi)$ (8)
@@
Therefore, the controller-plant model $\mathsf{M}_{\text{cpl}}$ can be constructed as a conservative non-probabilistic MDP (i.e., one whose behavior contains that of true system $\mathsf{M}_{\text{sys}}$) with a standard discrete abstraction technique (for example, in [10]). However, constructing a conservative $\textsf{M}_{\text{per}}$, either as an MDP or IMDP, is an open problem. Therefore, this paper focuses on synthesizing $\textsf{M}_{\text{per}}$ in a conservative manner using labeled perception data (i.e., the perception outputs $y$ paired with the true underlying states $x$) from the system $\mathsf{M}_{\text{sys}}$. In Section VI, we will show that our $\mathsf{M}_{\text{abs}}$ is a conservative system model, as per Equation 8.
@@
$\mathsf{M}_{\text{abs}}=\mathsf{M}_{\text{cpl}}~||~\textsf{M}_{\text{per}}$ (9)
@@
Our controller-plant model $\mathsf{M}_{\text{cpl}}$ is a non-probabilistic MDP with $\mathsf{M}_{\text{cpl}}=\{S_{\mathsf{M}_{\text{cpl}}},s^{0}_{\mathsf{M}_{\text{cpl}}},\alpha_{\mathsf{M}_{\text{cpl}}},\delta_{\mathsf{M}_{\text{cpl}}},L_{\mathsf{M}_{\text{cpl}}}\}$. The (discrete) state space $S$ is formed by partitioning the (continuous) state space $\mathcal{X}$ of $\mathsf{M}_{\text{sys}}$ into a finite set of mutually exclusive regions, so every $s\in S_{\mathsf{M}_{\text{cpl}}}$ has a corresponding region $X\subseteq\mathcal{X}$ (we use $x\in s$ to denote this correspondence). The initial state $s^{0}_{\mathsf{M}_{\text{cpl}}}$ is the state in $S_{\mathsf{M}_{\text{cpl}}}$ which contains the initial state $x_{0}$ of $\mathsf{M}_{\text{sys}}$ in its set of corresponding states. The actions $\alpha_{\mathsf{M}_{\text{cpl}}}=\alpha_{per}\times\alpha_{reach}$, where $\alpha_{per}=\{0,1\}$ (which we refer to as “perception actions”) are the outputs of the perception component $h$ and $\alpha_{reach}$ (which we refer to as “reachability actions”) are the actions that correspond to the non-determinism required to conservatively model the continuous system dynamics $f$ and controller $u$ with a discrete abstraction. [^2] For a visualization of this phenomenon, see Figure 2 in [10]. The transition relation $\delta_{\mathsf{M}_{\text{cpl}}}$ maps states and a pair (perception action, reachability actions) to the next state in $S_{\mathsf{M}_{\text{cpl}}}$. Finally, $L_{\mathsf{M}_{\text{cpl}}}:S\rightarrow 2^{AP}$, were $AP$ is a set of atomic propositions.
@@
$\mathsf{M}_{\text{cpl}}=\{S_{\mathsf{M}_{\text{cpl}}},s_{0},\alpha_{\mathsf{M}_{\text{cpl}}},\delta_{\mathsf{M}_{\text{cpl}}},L_{\mathsf{M}_{\text{cpl}}}\}$ is a conservative controller/plant abstraction if for every state $s\in S_{\mathsf{M}_{\text{cpl}}}$ and perception action $a_{per}\in\alpha_{per}$, the set of next states $S^{\prime}=\{s^{\prime}\in S\mid\exists a_{reach}\in\alpha_{reach}\text{ s.t. }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{a_{per},a_{reach}\},s^{\prime})=1\}$ in $\mathsf{M}_{\text{cpl}}$ contains every state in $\mathsf{M}_{\text{sys}}$ that can be mapped to by the dynamics of $\mathsf{M}_{\text{sys}}$. Formally:
$\forall s\in S_{\mathsf{M}_{\text{cpl}}},\forall a_{per}\in\alpha_{per},\forall x\in s,\exists s^{\prime}\in S,\exists a_{reach}\in\alpha_{reach}$
$\text{such that }\delta_{\mathsf{M}_{\text{cpl}}}(s,\{a_{per},a_{reach}\},s^{\prime})=1\text{ and }f(x,u(a))\in s^{\prime}$
@@
With the above in mind, this paper solves this problem:
@@
Given the closed-loop dynamical system $\mathsf{M}_{\text{sys}}$ defined in Equation 7, a safety property $\psi$, perception data $\mathcal{D}_{\text{sys}}=\{x_{i},z_{i}\}_{i=1,\ldots,N}$ where $x_{i}\in\mathbb{R}^{n}$ is the state of $\mathsf{M}_{\text{sys}}$ and $z_{i}\in\{0,1\}$ is the output of the perception component, and a conservative model $\mathsf{M}_{\text{cpl}}$ of the control and dynamics portion of $\mathsf{M}_{\text{sys}}$, synthesize a conservative model $\textsf{M}_{\text{per}}$ of the perception component(s) of $\mathsf{M}_{\text{sys}}$ such that
$P^{min}_{\mathsf{M}_{\text{cpl}}||\textsf{M}_{\text{per}}}(\psi)\leq P_{\mathsf{M}_{\text{sys}}}(\psi)$ (10)
V Conservative Approach to Perception Modeling
@@
    For each discrete state, compute an interval of detection probabilities using $\mathcal{D}$ and binomial confidence intervals [19]. This is a novel step where we differ from existing perception models [6, 11].
@@
V-A Step 1: State Space Discretization
@@
In terms of performing the binning in practice, we see two main ways of doing this. First, one can divide the state space into a set of equally spaced hyper-rectangles (or intervals if the state space is 1D, rectangles in 2D, …) [^3] This is the approach taken by the previous works [6, 11].. In this case, the size of the hyper-rectangle serves as a hyper-parameter of the approach. The second way is to first select the desired number of bins for each dimension of the state space $\mathcal{X}$, which we denote as $\mathcal{b_{1}},\ldots,\mathcal{b_{n}}$, and then divide the state space such that each bin has an equal number of datapoints in it. For example, if one has a 1D system and $l$ datapoints (so $|\mathcal{D}|=l$), then one would sort the points $\{x_{i},z_{i}\}\in\mathcal{D}$ by the $x_{i}$ values and assign each chunk of $\frac{\mathcal{l}}{n}$ to its own bin.
@@
We point out two other, subtler trade-offs at play that our discretization-based approach identifies. The existing perception abstraction approaches [6, 11] fail to account for these trade-offs, since they ignore the inherent uncertainty present in $\textsf{M}_{\text{per}}$— and so they fail to make it conservative. The first tradeoff is that using more bins means that there will be fewer datapoints per bin. This will lead to larger probability intervals for each bin in Step 2 since there are fewer datapoints fed into the binomial confidence intervals. These larger detection probability intervals in turn will lead to a larger interval on the estimated safety probability $\hat{P}(\psi)$. This trade-off captures the intuition that having fewer datapoints to compute the transition probabilities of $\textsf{M}_{\text{per}}$ generally results in a less precise model.
@@
Figure 2: Larger discretization bins lead to smaller binomial confidence intervals (blue) but necessitate more probability enlargement (orange).
V-B Step 2: Bin Probability Computation
@@
$\mathcal{D}_{s}=\{(x,z)\in\mathcal{D}~|~x\in s\}$ (11)
Just as in prior works [6, 11], we can then compute the empirical detection probability of the datapoints in $\mathcal{D}_{s}$:
$\hat{p}_{s}=\frac{\sum_{(x,z)\in\mathcal{D}_{s}}\mathbbm{1}_{z=1}}{|\mathcal{D}_{s}|}$ (12)
@@
To alleviate this issue, we replace the point estimate $\hat{p}_{s}$ with a confidence interval (CI), denoted as $[\underline{\hat{p}}_{s},\bar{\hat{p}}_{s}]$ for bin $s$ [^4] Using CIs to construct IMDPs has been explored in the areas of synthesis [26] and active learning [2] (which inspires our approach), but not verification. There are several methods to obtain a confidence interval for the detection probability of the perception model within the bin, also known as a binomial confidence interval. We employ the Clopper-Pearson confidence interval because its coverage level is never less than the desired nominal coverage [12]. When constructing a confidence interval, one needs to select a confidence level $1-\alpha$. Using a larger value of $1-\alpha$ leads to larger probability intervals, but the proportion of bins in which the confidence interval contains the true detection probability will, in turn, be larger.
[…]
$1-\sum_{s\in S}\frac{\alpha_{MC}}{|S|}=1-\alpha_{MC}$ (13)
@@
V-C Step 3: Probability Enlargement
@@
$\delta^{p}_{s}:=\left(\max_{x\in s}\;LR(x)\right)-\left(\min_{x\in s}\;LR(x)\right)$ (14)
@@
$[\underline{\hat{p}}_{s}-\delta^{p}_{s},\bar{\hat{p}}_{s}+\delta^{p}_{s}]$ (15)
@@
If we fit a surrogate model to detection probabilities, why not use it to construct $\textsf{M}_{\text{per}}$? We attempted to directly use the logistic regression, along with its associated confidence interval, to estimate the range of true detection probabilities within the bins. However, we found that the logistic regression confidence interval did not empirically achieve the coverage that we wanted. That is because the logistic regression CIs can be viewed as the CIs for the parameters of how to map data onto the space of logistic models — not for the data itself. We demonstrate this in the experiments in Section VII-D and particularly Figure 5(a).
VI Conservative Verification Guarantee
In this section, we state the main theoretical guarantee of this work. Specifically, we show that given the inputs to Problem 1 and an $\textsf{M}_{\text{per}}$ constructed using Steps 1–3 from Section V, verifying the resulting composed model $\mathsf{M}_{\text{abs}}$ will give a conservative estimate of the safety of the true system $\mathsf{M}_{\text{sys}}$:
$P_{\mathsf{M}_{\text{sys}}}(\psi)\in[Pr_{\mathsf{M}_{\text{abs}}}^{min}(\psi),Pr_{\mathsf{M}_{\text{abs}}}^{max}(\psi)]$ (16)